It is observed that the scores of a typical student in Math 477 on Exam I are normally distributed with mean 50 and standard deviation 5, while their scores on Exam II are independently and normally distributed with mean 55 and standard deviation 12. Suppose you are a typical student. What is the probability that the total of your scores on Exam I and II exceeds 150?
Answer #1

expected total score =50+55 =105

standard deviation =sqrt(52+122) =13

probability that the total of your scores on Exam I and II exceeds 150 =P(X>150)=P(Z>(150-105)/13)=P(Z>3.46)=0.0003